OpenStudy (anonymous):

5,941 mm =5.941 ? A.) m B.) km C.) cm D.) dm

OpenStudy (***[isuru]***):

10 mm = 1 cm so if u divide 5941 from 10 u will get the equivalent for cm 5941 mm = 5941/10 cm = 594.1 cm now 100 cm = 1m so if u divide 594.1 by 100 u will get the equivalent for m 5941 mm = 594.1 cm = 594.1/100 m = 5.941 m got it :) ?

OpenStudy (anonymous):

i wrote that down so I don't forget.

OpenStudy (***[isuru]***):

kk :P

OpenStudy (anonymous):

thanks for the help :)